# What Matters to Enhance Traffic Rule Compliance of Imitation Learning for Automated Driving

### v2

@@ Abstract
More research attention has recently been given to end-to-end autonomous driving technologies where the entire driving pipeline is replaced with a single neural network because of its simpler structure and faster inference time. Despite this appealing approach largely reducing the components in the driving pipeline, its simplicity also leads to interpretability problems and safety issues [1]. The trained policy is not always compliant with the traffic rules and it is also hard to discover the reason for the misbehavior because of the lack of intermediate outputs. Meanwhile, sensors are also critical to autonomous driving’s security and feasibility to perceive the surrounding environment under complex driving scenarios. In this paper, we proposed P-CSG, a penalty-based imitation learning approach with cross semantics generation sensor fusion technologies to increase the overall performance of end-to-end autonomous driving. In this method, we introduce three penalties - red light, stop sign, and curvature speed penalty to make the agent more sensitive to traffic rules. The proposed cross semantics generation helps to align the shared information from different input modalities. We assessed our model’s performance using the CARLA leaderboard - Town 05 Long benchmark and Longest6 Benchmark, achieving an impressive driving score improvement. Furthermore, we conducted robustness evaluations against adversarial attacks like FGSM and Dot attacks, revealing a substantial increase in robustness compared to baseline models. More detailed information, such as code base resources, and videos can be found at https://hk-zh.github.io/p-csg-plus.
Index Terms:
@@
Imitation Learning, Multi-modality Sensor Fusion, End-to-end Autonomous Driving
I Introduction
End-to-end autonomous driving integrates the perception and decision layers into one deep neural network. Enhancing the policy learning algorithm and improving perception abilities are two crucial aspects that can significantly enhance the overall performance of autonomous driving.
Imitation learning [2] and reinforcement learning [3] are two main learning paradigms unutilized end-to-end autonomous driving systems for the decision-making part. Although both approaches have been successful in the field of autonomous driving, they each have their own drawbacks. Imitation learning only learns the behaviors limited to the training data. This means that the model may struggle to generalize to new and unfamiliar situations. The quality of the expert policy also has a significant impact on the performance of imitation learning. The design of the reward function plays a critical role in the effectiveness of reinforcement learning for autonomous driving systems, and creating an appropriate reward function can be a complex task. Additionally, there is a risk that reinforcement learning models may learn unexpected and potentially dangerous behaviors in such safety-critical systems. In this article, we propose a penalty-based imitation learning paradigm that introduces the "penalty"/"reward" concept from reinforcement learing into imitation learning and aims to 1) achieve higher learning efficiency compared to reinforcement learning (learning from the expert demonstration) and 2) make the agent more sensitive to the traffic rules by involving some penalties designed in a simple manner.
The perception component in the neural network is used to extract essential information about the surrounding environment. Despite some Lidar-based approaches [4] [5], which leverages Lidar sensor input and HD maps, demonstrate impressive performance, relying on HD maps is not a feasible option as it requires substantial resources to create and maintain. They may not be universally available for all areas and regions. For this reason, an increasing number of research [6] [7] [8] [9] recently concentrate on multi-modality sensors (e.g Lidar and Camera) based approaches. Sensor fusion technologies are widely used in 2D object and 3D object detection. Classified by the types of point cloud representation of Lidar input, there exist three main approaches, namely point-based [10] [11], voxel-based [12] [13] [14] [15], and range-view-based [16] [17] [18] [19] approaches. More recently, with the popularity of attention mechanisms, many researchers are trying to use Transformer [20] to integrate multimodal information [7] [21] [8].In this paper, We shift our approach by extracting and aligning shared information from diverse modalities, enabling the model to extract the global context within multi-modalities
@@ I Introduction
We propose Penalty-based Imitation Learning with Cross Semantics Generation (P-CSG). We evaluate our model on Carla Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark. It achieves state-of-the-art performance, especially in collision and traffic rule violations. We also analyze the performance and robustness of our autonomous driving model under two types of adversarial attacks: FGSM and Dot attacks. The results show that our model outperforms the other baselines from the perspective of safety.
@@
  • Our innovative multi-sensor fusion technique involves aligning shared information from various modalities, thus simplifying the process of extracting the global context across diverse modalities.
@@
II Related Works
There are currently two primary methods used in autonomous driving technology: modular and end-to-end approaches. The modular approach involves a detailed series of software modules that work together to control the vehicle. By employing a modular approach, engineering teams can focus on specific and well-defined subtasks and make enhancements independently throughout the entire system. This allows the system to remain operational as long as the intermediate outputs remain functional. On the other hand, the end-to-end approach considers the entire driving process as one comprehensive learning task. It leverages the ability of the neural network to learn task-specific features and perform task-specific actions. Contrary to the modular approach, the agent learns without relying on explicit rules or hand-engineered features, namely the human-defined information bottlenecks. This means that the vehicle can adapt to different driving scenarios and environments and potentially achieve better performance than systems that rely on traditional rule-based approaches. End-to-end approaches have shown great success in computer vision, such as object detection, object tracking, and semantic segmentation. The success of these tasks builds a solid foundation for end-to-end autonomous driving. It is reasonable to assume that in the future, the end-to-end approach will also have the ability to operate a vehicle autonomously. Lately, there has been a growing trend in research to utilize behavior cloning for training agents to accomplish autonomous driving tasks. This involves the agent learning from demonstrations provided by experts in the field [7] [8] [6] [9]. They collect the demonstration in CARLA [22], a simulation urban environment in which the researchers and developers can test and evaluate their algorithms and models for autonomous driving. They employ strategies like enhancing multi-modality sensor fusion [7] [8], introducing safety module [8] to avoid dangerous behavior and improve the overall performance of autonomous driving. In our approach, penalties for traffic rules are introduced to ensure the safety of driving. We also introduce the cross semantics generation, which generates the semantic information of the other modality, to capture the global context across different modalities.
III Methodologies
@@
Feature ExtractionCross Semantic GenerationAuxiliaryPenalty-based Imitation LearningRefer to captionRefer to captionRefer to captionptResnet50ptResnet18pt512pt512ptDecoderRefer to captionRefer to captionptMLPptMLPContractive Losspt128Refer to captionRefer to captionRefer to captionpt4Measurementspt1156Refer to captionRed Light PenaltyRefer to captionStop Sign PenaltyRefer to captionSpeed PenaltyptMLPpt64ptGRUptptGoalLocationptptptWaypointsptPIDRefer to captionRefer to captionPenalties
Figure 1: This is an overview of our proposed Penalty-based Imitation Learning with Cross Semantics Generation. Let’s start with the first feature extraction part. The top-down LiDAR pseudo image is fed into a ResNet18 architecture to extract the features from the LiDAR input. The front camera image is fed into a ResNet50 architecture to extract the features from the camera input. Then, the Cross Semantic Generation is conducted to cross-generate the front and top-down semantic segmentations in the frame of the front image and LiDAR pseudo image. We apply the contrastive loss the align the shared features of the camera and LiDAR inputs. The resampled shared embedding from latent space will be used to generate the waypoints. The features extracted from the front camera image are also leveraged to perform auxiliary tasks ’stop sign indication’ and ’red light indication’. In terms of penalty-based imitation learning, the LiDAR features, camera features, shared features, and measurements are concatenated and fed into an MLP network to reduce the dimension. A GRU structure is followed to recurrently generate the waypoints, also taking the goal location as input. The generated waypoints will be applied with the three penalties, namely, red light penalty, stop sign penalty, and speed penalty. The task of the PID controller is to map the waypoints into the action of steer, throttle, and brake.
III-A Problem Setting
The task we concentrate on is point-to-point navigation in an urban setting where the goal is to complete a route with safe reactions to dynamic agents such as moving vehicles and pedestrians. The traffic rules like red lights and stop signs should be followed. We consider the problem of learning a goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$ that outputs action $\bm{a}\in\mathcal{A}$, conditioned on the current observed state $\bm{s}\in\mathcal{S}$ and a high-level goal location $\bm{g}$ provided by GPS. The environment we used to train and test the performance of the trained agent is CARLA. The specific environment setting we used can be characterized by the following statements:
  • A state space $\mathcal{S}\subset\{(\mathbb{R}^{H_{\textrm{c}}\times W_{\textrm{c}}\times 3},\mathbb{R}^{H_{\textrm{L}}\times W_{\textrm{L}}\times 2},\mathbb{R}^{4})\}$ which is the combination of the observation of camera with the shape of $H_{\textrm{c}}\times W_{\textrm{c}}\times 3$, bird’s eye view LiDAR input with the shape of $H_{\textrm{L}}\times W_{\textrm{L}}\times 2$, and measurements of current throttle, brake, steer, and velocity.
  • A multidimensional action space $\mathcal{A}\subset\mathbb{R}^{3}$. The action space contains the parameters of steer, throttle, and brake to drive the agent to finish tasks.
@@ III-A Problem Setting
III-A1 Observation Space
Similar to prior works [23], [24] [7] [21], the LiDAR point cloud data is transformed into a 2-bin histogram by mapping it onto a 2D BEV grid of fixed resolution. To encompass a 32m x 32m BEV grid, we limit our analysis to points located within 32m in front of the ego vehicle and 16m on each side. We partition the grid into blocks of 0.125m $\times$ 0.125m, which gives us a resolution of 256 $\times$ 256 pixels. The height dimension is discretized into two bins, distinguishing points that are either on/below or above the ground plane. This results in a two-channel pseudo image input $\mathbb{R}^{256\times 256\times 2}$.
In terms of RGB input, we leverage three cameras positioned to face forward, 60 degrees to the left, and 60 degrees to the right. These cameras have a horizontal field of view (FOV) of 120 degrees. The image from the cameras has a resolution of $400\times 300$ pixels. The forward image is cropped to $300\times 160$ pixels and the 60 degrees to the left and right images are cropped to $160\times 234$ pixels. we want to use cropping to eliminate radial distortion at the edges. We combine these three cropped images into a single image input with a resolution of $768\times 160$ pixels.
In addition, the current throttle, brake, steering, and velocity measurements are included in the observation space to provide a point of reference for the agent.
III-A2 Action Space
Instead of directly predicting the next action of throttle, steer, and brake, we estimate the future waypoints $\mathcal{W}$ of the ego-vehicle in bird’s eye view space which is centered at the ego vehicle’s current coordinate frame. The trajectory consists of a sequence of 2D waypoints $\mathcal{W}=\{\bm{w_{t}}=(x_{t},y_{t})\}^{T}_{t=1}$. In our settings, $T=4$ which means the agent needs to predict 4 waypoints in the future for each frame. Then we use two fine-tuned PID controllers for lateral and longitudinal control to obtain steer, throttle, and brake values from the predicted waypoints. These two PID controllers are not involved in the training process.
III-A3 Goal Location
According to CARLA [22] 0.9.10’s protocol, the high-level goal locations $\bm{g}$ are provided as GPS coordinates. This goal location $\bm{g}$ is sparse (hundreds of meters apart) which can only be used as guidance. We kindly remind the readers that the goal location is different from those predicted waypoints, which are dense, and only a few meters away from each other.
III-B Policy Learning
@@
In this paper, we plan to use the imitation learning mechanism to learn the goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$. As we discussed previously, we intend to predict the future waypoints instead of directly learning the actions. An expert policy is applied in the environment to collect a dataset $\mathcal{D}=\{(\bm{s},\mathcal{W})\}^{N}$ with the size of $N$, which contains multiple tuples of the current observation $\bm{s}$ and a set of waypoints $\mathcal{W}$ in the future timesteps. Similar to previous works [7], [25], we also use $l_{1}$-norm based loss function because of its demonstrated robustness to outliers. For each input, the policy learning loss can be formalized as:
$\mathcal{L}_{\textrm{pl}}=\sum_{t=1}^{T}||\bm{\hat{w}_{t}}-\bm{w_{t}}||_{1}$ (1)
@@
where $\bm{\hat{w}_{t}}$ is the $t$-th predicted waypoint and $\bm{w_{t}}$ is the t-th ground truth waypoint produced by the expert policy. $\mathcal{L}_{\textrm{pl}}$ is the loss function of policy learning.
III-C Cross Semantics Generation
@@
The motivation for our approach is based on the fact that multi-modality inputs have common information and also their own unique information. For instance, the vehicle’s and pedestrian’s shape and location are the shared information of LiDAR and Camera input. The term unique information refers to information that is not present in other sources of input. Extracting and aligning the common information from both modalities, helps the decision network to find the logical connection between modalities and capture their global context since the extracted features are more reasonable and easier to understand. There are two problems we want to solve in this section. They are 1) How to find the shared information of these two modalities? 2) How to align the extracted shared information in the same space?
@@
To address the first problem, we propose the cross semantics generation approach. We plan to utilize the LiDAR input to produce semantic information from the camera and also use the camera input to generate semantic information from the LiDAR. As Figure 1 demonstrates, the information flows for the LiDAR and camera semantic information generation are crossed since we used the information from one modality to generate the semantic segmentation of the other modality. That is the reason we name our approach cross semantics generation. The RGB image and the bird’s eye view of the LiDAR pseudo image are fed into two residual networks [26] to extract the features of these two modalities. We utilize two linear layers to extract the shared information of the camera and LiDAR. Those from the camera input extracted features are used to generate the top-down semantic segmentation which is aligned with the LiDAR input and those from the LiDAR input extracted features are used to generate the semantic segmentation of the front image. The reason we used the information to generate the semantic segmentation is that 1) semantic segmentation has less noise compared to the original data and 2) with the semantic segmentation we can filter out the unique information such as traffic light color and traffic sign pattern. In our setup, the semantic segmentation contains 4 channels, namely the drivable area, the non-drivable area, the object in the drivable areas (vehicle and pedestrians), and others.
@@
We define $y_{\textrm{front}}$ as the ground truth front segmenation tensor with the shape of $160\times 768\times 4$ and $\hat{y}_{\textrm{front}}$ as the output of the front view decoder with the same shape. We also define $y_{\textrm{td}}$ as the ground truth top-down segmentation tensor with the shape of $256\times 256\times 4$ and $\hat{y}_{\textrm{td}}$ as the output of the top-down view decoder with the same shape. The reconstruction losses can be written as $\mathcal{L}_{\textrm{front}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{front}},y_{\textrm{front}})$ and $\mathcal{L}_{\textrm{td}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{td}},y_{\textrm{td}})$, where $\mathcal{L}_{\textrm{CE}}(\cdot)$ indicates the cross entropy loss. $\mathcal{L}_{\textrm{front}}$ and $\mathcal{L}_{\textrm{td}}$ represent the loss function of front reconstruction loss and top-down reconstruction loss, respectively.
@@
To align the extracted shared information into the same space, we leverage the contrastive loss. The primary goal of contrastive loss is to learn an embedding space where similar examples are placed closer together, while dissimilar examples are placed farther apart. This is achieved by defining a distance metric between the embeddings and then minimizing the distance between similar examples while maximizing the distance between dissimilar examples. Note that the embeddings extracted from LiDAR and camera inputs are mapped into two Gaussian distributions by generating mean and variance, using MLP networks. Suppose $P_{1}$ and $P_{2}$ represent two batches of Gaussian distributions, derived from RGB and LiDAR inputs. Each batch contains $N_{b}$ number of distributions. The contrastive loss can be formulated as follows:
$\mathcal{L}_{\textrm{align}}({P_{1}},{P_{2}})=\bm{E}\odot\bm{D}+(1-\bm{E})\odot\max\{0,\epsilon_{\textrm{a}}-\bm{D}\}$ (2)
@@
where $\bm{D}$ is the matrix consisting of $d_{ij}=d(p_{1}^{i},p_{2}^{j})$. $p_{1}^{i}$ is the $i$-th distribution in ${P_{1}}$ and $p_{2}^{j}$ is the $j$-th distribution in ${P_{2}}$. $d(p_{1}^{i},p_{2}^{j})$ indicates the metric to measure the difference of two distributions. $\bm{E}$ is the matrix consisting of $e_{ij}$ which indicates if $p_{1}^{i}$ and $p_{2}^{j}$ belong to the same frame. $\epsilon_{\textrm{a}}$ is the threshold difference between similar samples and dissimilar samples. $\odot$ represents the Hadamard product of two matrices. In our setting, we use the symmetric version of Kullback–Leibler divergence
$d(p_{1}^{i},p_{2}^{j})=\frac{1}{2}\textrm{KL}(p_{1}^{i}||p_{2}^{j})+\frac{1}{2}\textrm{KL}(p_{2}^{j}||p_{1}^{i})$ (3)
@@
where $\textrm{KL}(\cdot)$ is the KL-divergence metric, as the metric to measure the difference between two distributions. The sampled shared embedding is concatenated with image and LiDAR embeddings, along with measurements (velocity, throttle, steer, and brake from the previous frame). This combined information serves as the input for the penalty-based imitation learning module.
III-D Auxiliary Tasks

### v3

@@ Abstract
End-to-end autonomous driving, where the entire driving pipeline is replaced with a single neural network, has recently gained research attention because of its simpler structure and faster inference time. Despite this appealing approach largely reducing the complexity in the driving pipeline, it also leads to safety issues because the trained policy is not always compliant with the traffic rules. In this paper, we proposed P-CSG, a penalty-based imitation learning approach with contrastive-based cross semantics generation sensor fusion technologies to increase the overall performance of end-to-end autonomous driving. In this method, we introduce three penalties - red light, stop sign, and curvature speed penalty to make the agent more sensitive to traffic rules. The proposed cross semantics generation helps to align the shared information of different input modalities. We assessed our model’s performance using the CARLA Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark, achieving 8.5% and 2.0% driving score improvement compared to the baselines. Furthermore, we conducted robustness evaluations against adversarial attacks like FGSM and Dot attacks, revealing a substantial increase in robustness compared to other baseline models.
Keywords:
@@
Imitation Learning Multi-modality Sensor Fusion End-to-end Autonomous Driving
1 Introduction
End-to-end autonomous driving [20] integrates the perception and decision-making layers into one deep neural network. The perception component extracts essential information about the surrounding environment. Despite some approaches [4] [32], which leverage LiDAR sensor input and HD maps and demonstrate impressive performance, relying on HD maps is not a feasible option as it requires substantial resources to create and maintain. They may not be universally available for all areas and regions. Recent research has concentrated on multi-modality sensor techniques utilizing both LiDAR [24, 15] and cameras [35, 12, 33, 6]. To manage the complexity of LiDAR input, various approaches have been developed, including point-based [44], voxel-based [34, 46, 48, 41], and range-view-based approaches [14, 8, 36].
More recently, with the popularity of attention mechanisms, many researchers are trying to use Transformer [39, 43] to integrate multimodal information [12] [29] [33]. Despite its potential, the Transformer architecture on a large scale leads to a significant increase in both the training and inference time. In this paper, we shift our approach by extracting and aligning shared information from diverse modalities by contrastive learning, enabling the model to extract the global context within multi-modalities.
Refer to caption(a)(b)
Figure 1: (a) Common imitation learning approach could learn the wrong behavior which violates the traffic rule. (b) Our approach penalizes the predicted waypoints violating traffic rules to enhance the agent’s traffic rule adherence.
Imitation learning [1] and reinforcement learning [37] are two main learning paradigms utilized by end-to-end autonomous driving systems for the decision-making part. Reinforcement learning suffers from the low sample efficiency and the complexity of reward signs. It may also learn risky actions, leading to accidents and unsafe driving. Due to these shortages, other researchers have turned to imitation learning approaches [39, 12, 29, 33]. However, these approaches still lack effective mechanisms to ensure the agent to comply the traffic rules. For one thing, the expert policy to generate demonstrations may still make mistakes when collecting the training data. Also, the metrics for autonomous driving (e.g., traffic rule violations) and the objectives of imitation learning (minimize the difference between the predicted and ground truth waypoints) remain disparate, indicating that a low loss in the learning objective does not guarantee optimal performance of the agent in the testing environments.
To overcome the above-mentioned limitations, we propose Penalty-based Imitation Learning with contrastive-based Cross Semantics Generation (P-CSG). Our approaches penalize the behavior (higher loss) that violates the traffic rules during the imitation learning process to make the agent better adhere to these traffic rules (cf. Figure 1). Meanwhile, we adopt a contrastive-based sensor fusion technology that aligns the shared information of multi-modalities so that the model can better understand the global context within multi-modalities.
@@ 1 Introduction
We evaluated our model on the CARLA Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark. It achieves outstanding performance, especially in collision and traffic rule violations. We also analyze the performance and robustness of our autonomous driving model under two types of adversarial attacks: Fast Gradient Sign Method (FGSM) Attack [17] and Dot attacks [22]. The results show that our model outperforms the other baselines from the safety perspective.
@@
  • Our innovative multi-sensor fusion technique involves aligning shared information from various modalities, thus extracting the global context across diverse modalities.
@@
  • We evaluate our method’s performance under adversarial attacks w.r.t. FSGM and Dot Attacks. The results show that our model adopts a more cautious strategy and has a more robust performance.
The rest of this paper is organized as follows: Section 2 offers a detailed review of related literature. We present our methodology in Section 3. Section 4 provides the experiment results and discussions. Section 5 conduct robustness study against adversarial attacks. We conclude our paper in Section 6.
2 Related Works
Imitation learning [35, 29, 6, 12, 33, 2] are widely used in the field of autonomous driving [9, 10, 11, 38]. Many approaches use the CARLA simulator [13] to collect training data and test the performance of the trained agent in the simulation environment. To improve the agent’s overall performance, many approaches investigate the multi-sensor fusion technologies and safety mechanisms.
2.1 Multi-sensor Fusion Technologies
Multi-sensor fusion has received much research attention in the field of end-to-end autonomous driving. Recent works [42, 3, 47, 28], complementing camera images with depth and semantics has shown the potential to improve the overall driving performance. More research attention has been given to LiDAR and camera input fusion since they are complementary to each other in terms of representing the scene and are readily available in autonomous driving systems. Early works such as LateFusion [35] concatenate the feature embeddings from different modalities and use a Multi-Layer-Perception (MLP) network to weigh, select, and fuse the concatenated features. TransFuser [29] leverage Transformer-based approaches to achieve multi-modal global context. They apply cross-attention to the feature embeddings from different modalities in different layers. TransFuser+ [12], as an extension of TransFuser, introduces auxiliary tasks: depth prediction and semantic segmentation are performed with image information, while HD map prediction and vehicle object detection are handled with bird’s eye view (BEV) LiDAR information. These auxiliary tasks serve to illuminate the inner workings of the entire network. Similarly, InterFuser [33] also adopts the Transformer-based sensor fusion approach.
Contrastive Learning, which is widely used in vision language models (VLMs) [30, 23], demonstrates its strong performance in multi-modal information handling. Some recent works [26, 45, 25] focus on using contrastive learning for multi-modal information fusion. In this paper, we investigate contrastive learning for LiDAR and camera input fusion and propose a contrastive-based cross semantic generation sensor fusion method to extract the global context within the modalities of LiDAR and camera.
2.2 Safety Mechanism
To bolster the safety of autonomous driving agents, researchers in [12] concentrate on refining the expertise of the expert agent. By filtering out incorrectly demonstrated actions executed by the expert policy, the agent avoids being misled and can more effectively acquire skills from the demonstrations. LAV [6] supervises waypoint outputs with additional data by making predictions for other nearby agents. InterFuser [33] developed a safety control module to regulate the agent’s behaviors, preventing the agent from violating the traffic rules. However, this control module heavily depends on hand-designed heuristics and can vary in different systems. In our paper, we introduce the ‘penalty’ concept to the IL framework, which incentivizes the trained agent to adopt safer driving behaviors in an end-to-end manner.
2.3 Combining IL and RL
Methods such as DQfD[19], DDPGfD[40], and DAPG[31] have shown that integrating IL with RL can help address exploration challenges in domains with sparse rewards. Other offline RL approaches, such as TD3+BC [16] and CQL [21], combine RL and IL objectives to regularize Q-learning updates and prevent overestimation of out-of-distribution values. In the autonomous driving field, Lu et al. [27] proposed behavior cloned soft actor-critic (BC-SAC), which incorporates a behavioral cloning objective into the SAC objective.
Our approach starts with an IL perspective by directly incorporating the concept of penalties and rewards from RL into IL. We model penalties as constraints and violating these constraints leads to higher loss. Q-function learning is not included in our approach.
3 Methodologies
@@
3.1 Problem Setting
We concentrate on point-to-point navigation in an urban setting where the goal is to complete a route with safe reactions to dynamic agents such as moving vehicles and pedestrians. Traffic rules should be followed, such as red lights and stop signs. We consider the problem of learning a goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$ that outputs action $\bm{a}\in\mathcal{A}$, conditioned on the current observed state $\bm{s}\in\mathcal{S}$ and a goal location $\bm{g}$ provided by GPS. The environment we used to train and test the performance of the trained agent is CARLA. The specific environment setting we used can be characterized by the following statements:
  • A state space $\mathcal{S}\subset\{(\mathbb{R}^{H_{\textrm{c}}\times W_{\textrm{c}}\times 3},\mathbb{R}^{H_{\textrm{L}}\times W_{\textrm{L}}\times 2},\mathbb{R}^{4})\}$ which is the combination of the observation of camera with the shape of $H_{\textrm{c}}\times W_{\textrm{c}}\times 3$, bird’s eye view (BEV) LiDAR pseudo image input with the shape of $H_{\textrm{L}}\times W_{\textrm{L}}\times 2$, and measurements of current throttle, brake, steer, and speed.
  • A multidimensional action space $\mathcal{A}\subset\mathbb{R}^{3}$. The action space contains the parameters of steer, throttle, and brake to drive the agent to finish tasks. Instead of directly predicting the next action of throttle, steer, and brake, we estimate the future waypoints $\mathcal{W}$ of the ego-vehicle in BEV space, which is centered at the ego vehicle’s current coordinate frame. Then, we use two fine-tuned PID controllers for lateral and longitudinal control to obtain steer, throttle, and brake values from the predicted waypoints.
@@ 3.1 Problem Setting
3.2 Policy Learning
@@
In this paper, we use the imitation learning mechanism to learn the goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$. As we discussed previously, the future waypoints instead of the actions are predicted. An expert policy is applied in the environment to collect a dataset $\mathcal{D}=\{(\bm{s},\mathcal{W})\}^{N}$ with the size of $N$, which contains tuples of the observation $\bm{s}$ and a set of waypoints $\mathcal{W}$ in future time steps. Similar to previous works [12], [7], we use $l_{1}$-norm based loss function because of its demonstrated robustness to outliers. For each input, the policy learning loss can be formalized as:
$\mathcal{L}_{\textrm{pl}}=\sum_{t=1}^{T}||\bm{\hat{w}_{t}}-\bm{w_{t}}||_{1}\text{,}$ (1)
@@
where $\bm{\hat{w}_{t}}$ is the $t$-th predicted waypoint and $\bm{w_{t}}$ is the t-th ground truth waypoint produced by the expert policy.
[…]
Figure 2: An Overview of Our Penalty-based Imitation Learning with Cross Semantics Generation.
3.3 Contrastive-based Cross Semantics Generation
@@
Our approach’s motivation is based on the fact that multi-modality inputs have common information and also their own unique information. For instance, the vehicle’s and pedestrian’s shape and location are the shared information of LiDAR and camera input. Unique information refers to information that is not present in the other source of modality. Extracting and aligning the common information from both modalities can capture the global context within multi-modalities. In this section, we address two problems. 1) How to find the shared information of these two modalities? 2) How to align the extracted shared information in the same space?
@@
To address the first problem, we propose the cross semantics generation approach. We utilize the LiDAR input to produce semantic segmentation of the camera input and use the camera input to generate semantic segmentation of the LiDAR pseudo image input. As Figure 2 demonstrates, the information flows of the LiDAR (gray) and camera (black) semantic information generation are crossed since we used the information from one modality to generate the semantic segmentation of the other modality. This is the reason that we name our approach cross semantics generation. The RGB image and the BEV of the LiDAR pseudo image are fed into two residual networks [18] to extract the features of these two modalities. We utilize two linear layers to extract the information shared by the camera and LiDAR. The feature embeddings are leveraged to generate the semantic segmentation because 1) semantic segmentation has less noise than the original data, and 2) with semantic segmentation, we filter out the unique information, such as traffic light color and traffic sign patterns. In our setup, the semantic segmentation contains 4 channels: the drivable area, the non-drivable area, the objects (vehicles and pedestrians) in the drivable areas, and others.
@@
We define $y_{\textrm{front}}$ as the ground truth front segmenation tensor with the shape of $160\times 768\times 4$ and $\hat{y}_{\textrm{front}}$ as the output of the front view decoder with the same shape. We also define $y_{\textrm{td}}$ as the ground truth top-down segmentation tensor with the shape of $256\times 256\times 4$ and $\hat{y}_{\textrm{td}}$ as the output of the top-down view decoder with the same shape. The reconstruction losses can be written as $\mathcal{L}_{\textrm{front}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{front}},y_{\textrm{front}})$ and $\mathcal{L}_{\textrm{td}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{td}},y_{\textrm{td}})$, where $\mathcal{L}_{\textrm{CE}}(\cdot)$ indicates the cross entropy loss. $\mathcal{L}_{\textrm{front}}$ and $\mathcal{L}_{\textrm{td}}$ represent the functions of front reconstruction loss and top-down reconstruction loss, respectively.
@@
To address the second problem, we leverage the contrastive loss to align the extracted shared information into the same space. The primary goal of contrastive loss is to learn an embedding space where similar examples are placed closer together while dissimilar examples are placed farther apart. This is achieved by defining a distance metric between the embeddings and then minimizing the distance between similar examples while maximizing the distance between dissimilar examples. The embeddings extracted from LiDAR and camera inputs are mapped into two Gaussian distributions by generating mean and variance using MLP networks. Suppose $P_{1}$ and $P_{2}$ represent two batches of Gaussian distributions derived from RGB and LiDAR inputs. Each batch contains $N_{b}$ number of distributions. The contrastive loss can be formulated as follows:
$\mathcal{L}_{\textrm{align}}({P_{1}},{P_{2}})=mean(\bm{E}\odot\bm{D}+(1-\bm{E})\odot\max\{0,\epsilon_{\textrm{a}}-\bm{D}\})$ (2)
@@
where $\bm{D}$ is the matrix consisting of $d_{ij}=d(p_{1}^{i},p_{2}^{j})$. $p_{1}^{i}$ is the $i$-th distribution in ${P_{1}}$ and $p_{2}^{j}$ is the $j$-th distribution in ${P_{2}}$. $d(p_{1}^{i},p_{2}^{j})$ indicates the metric to measure the difference of two distributions. $\bm{E}$ is the matrix consisting of $e_{ij}$ which indicates if $p_{1}^{i}$ and $p_{2}^{j}$ belong to the same frame. $\epsilon_{\textrm{a}}$ is the threshold difference between similar samples and dissimilar samples. $\odot$ represents the Hadamard product of two matrices. $mean(\cdot)$ represents the mean value of one tensor. In our setting, we use the symmetric version of Kullback–Leibler divergence to measure the distance of two distributions:
$d(p_{1}^{i},p_{2}^{j})=\frac{1}{2}\textrm{KL}(p_{1}^{i}||p_{2}^{j})+\frac{1}{2}\textrm{KL}(p_{2}^{j}||p_{1}^{i})\text{,}$ (3)
@@
where $\textrm{KL}(\cdot)$ is the KL-divergence metric as the metric to measure the difference between two distributions. The sampled shared embedding is concatenated with image and LiDAR embeddings, along with measurements (speed, throttle, steer, and brake from the previous frame). This combined information serves as the input for the penalty-based imitation learning module.
3.4 Auxiliary Tasks